Headland Inc. manufactures cycling equipment. Recently, the vice president of operations of the company has requested construction of a new plant to meet the increasing demand for the company’s bikes. After a careful evaluation of the request, the board of directors has decided to raise funds for the new plant by issuing $3,078,200 of 12% term corporate bonds on March 1, 2020, due on March 1, 2035, with interest payable each March 1 and September 1, with the first interest payment on September 1st, 2020. At the time of issuance, the market interest rate for similar financial instruments is 10%.

As the controller of the company, determine the selling price of the bonds.

Selling Price of the bonds?___________
Answer #1

Selling price of bonds is equal to the present value of all future coupon payments and the principal amount

Semi-annual required return = 10%/2 = 5%

Number of semi-annual periods = 15*2 = 30 periods

Hence, Selling price of bonds = 3,078,200*12%*1/2*PVAF(5%, 30 periods) + 3,078,200*PVF(5%, 30 periods)

= 184,692*15.3725 + 3,078,200*0.2314

= $3,551,473.25

Hence, selling price of bonds = $3,551,473.25